One number is 8 less than another. If 9 times the smaller number minus 4 times the larger number is 8, find the two numbers.
Answer by josmiceli(19441)

x = smaller number
y = larger number
x = y - 8
9x - 4y = 8
subtract y from both sides of the 1st equation
x - y = -8
9x - 4y = 8
Add the equations
10x -5y = 0
Add 5y to both sides
10x = 5y
divide both sides by 5
2x = y
substitute 2x for y in 1st equation
x - 2x = -8
-x = -8
x = 8 answer
y = 2x
y = 16 answer
check
9x - 4y = 8
9*8 - 4*16 = 8
72 - 64 = 8
8 = 8
OK